Calculate the number of atoms in a 7.84×103 g sample of cesium.
1
Expert's answer
2020-08-05T05:10:47-0400

1 mole of any compound contains 6.022 * 1023 atoms.

1 mole of cesium weighs 133 g.

Find how many moles of cesium do we have:

"\nu" (Cs) = 7.84 * 10^3 / 133 = 58.95 (mol)

Find how many atoms of cesium do we have:

N(Cs) = 58.95 * 6.022 * 1023 = 3.55 * 1025 (atoms)


Answer: there are 3.55 * 1025 atoms in a 7.84×103 g sample of cesium.